OpenStudy (anonymous):

Find the products of each (x + 4)(x − 4) and (x + 4)(x + 4)

OpenStudy (saifoo.khan):

(x + 4)(x − 4) x^2 - 4x + 4x - 16 x^2 -16

jimthompson5910 (jim_thompson5910):

(x + 4)(x − 4) x^2-4x+4x-16 x^2-16 So (x + 4)(x − 4) = x^2-16 ----------------------------- (x + 4)(x + 4) x^2+4x+4x+16 x^2+8x+16 So (x + 4)(x + 4) = x^2+8x+16

OpenStudy (anonymous):

(x + 4)(x − 4)=x^2-16 (x + 4)(x + 4)=x^2+8x+16

OpenStudy (saifoo.khan):

Right nick.